A spacecraft is in orbit at an altitude of 100 km above the surface of the moon.




What is its speed? Mmoon = 7.34 x 1022 kg and Rmoon = 1.84 x 106 m





1
Expert's answer
2022-06-03T15:25:28-0400
"v=\sqrt{G\frac{M}{r}}\\
=\sqrt{6.67*10^{-11}*\frac{7.34*10^{22}}{1.84*10^6+100*10^3}}=1589\:\rm m/s"